# Exploring Test-Time Adaptation for Object Detection in Continually Changing Environments

### v3

@@ C.3. Additional Ablation Study
Table 6. Additional Ablation Study. The “CTAOD-unstop” denotes CTAOD without dynamic skipping. The “FT” represents replacing dynamic threshold in CTAOD with a fixed threshold. The “SR” denotes the stochastic restoration (Wang et al., 2022), and the “DR” indicates the data-driven restoration (Brahma and Rai, 2023). All experiments are done on long-term Cityscapes-to-Cityscapes-C tasks.
As presented in Table 6, we perform an additional ablation study to evaluate the effectiveness of the proposed Dynamic Threshold (DT) strategy in Adaptive Monitoring and Data-driven Stochastic Restoration (DSR). The findings indicate that the DT approach results in improvements of 1.3, 0.8, and 0.5 mAP over fixed thresholds set at 0.9, 0.8, and 0.7, respectively. While the stochastic restoration and data-driven restoration limit the performance to 27.1 and 28.3 mAP, the DSR mechanism raises the performance to 29.0 mAP.
@@ Appendix E Limitation
Firstly, our experiments are primarily based on a single detector backbone, i.e. Faster R-CNN (Ren et al., 2015). Future work will extend to other advanced detector backbone networks to assess the universality and generalization capabilities of CTAOD. Moreover, the proposed DSR mechanism to reset the model every iteration increases the computational cost. Future work could explore more efficient mechanisms for restoration, such as restoring the model only when significant changes in the distribution of the target domain are detected. Finally, the evaluation tasks in our work are designed to simulate real-world adaptation scenarios by incorporating datasets affected by corruption and adverse conditions. However, real-world data distributions are inherently more complex. Consequently, a promising avenue for future research would be to apply our methodology in practical, real-world systems to further validate its effectiveness.
@@
26:       // 3. Data-driven Stochastic Restoration
@@
(f) CTAOD (ours)
Figure 4. Complete qualitative results. We compare the detection results of the CTAOD and other baseline methods in the $10^{th}$ round of adaption to Brightness corruption on the long-term Cityscapes-to-Cityscapes-C task.
@@
(f) CTAOD (ours)
Figure 5. Complete qualitative results. We compare the detection results of the CTAOD and other baseline methods in the $10^{th}$ round of adaption to Defocus corruption on the long-term Cityscapes-to-Cityscapes-C task.
@@
(f) CTAOD (ours)
Figure 6. Complete qualitative results. We compare the detection results of the CTAOD and other baseline methods in the $10^{th}$ round of adaption to Fog corruption on the long-term Cityscapes-to-Cityscapes-C task.
@@
(f) CTAOD (ours)
Figure 7. Complete qualitative results. We compare the detection results of the CTAOD and other baseline methods in the $10^{th}$ round of adaption to Motion corruption on the long-term Cityscapes-to-Cityscapes-C task.
@@
(f) CTAOD (ours)
Figure 8. Complete qualitative results. We compare the detection results of the CTAOD and other baseline methods in the $10^{th}$ round of adaption to Snow corruption on the long-term Cityscapes-to-Cityscapes-C task.

### v4

@@ 4.4.3. Qualitative Results
Table 6. Fine-grained ablation study. The “FT” represents replacing AM module in AMROD with a fixed threshold. The “SR” and the “DR” indicate replacing ARR with the stochastic restoration (Wang et al., 2022) and the data-driven restoration (Brahma and Rai, 2023), respectively. All experiments are done on long-term Cityscapes-to-Cityscapes-C tasks.
4.4.4. Ablation Study
We conduct an ablation study to evaluate the impact of the OCL, AM, and ARR. As shown in Table 5, integrating the mean-teacher architecture achieves the performance of 25.9 mAP. Building upon this foundation, the proposed OCL, AM, and ARR achieve an additional 1.5, 1.4, and 0.8 mAP improvement individually. Moreover, the combination of the OCL and AM modules raises the performance to 28.6 mAP. Furthermore, the inclusion of the ARR mechanism brings a 0.6 mAP boost in performance, culminating in a comprehensive model performance of 29.2 mAP.
As presented in Table 6, we further perform a fine-grained comparison study to assess the effectiveness of the AM and ARR modules. This is done by replacing AM and ARR with a Fixed threshold (FT), stochastic restoration (SR) (Wang et al., 2022), or data-driven restoration (DR) (Brahma and Rai, 2023). The findings indicate that the dynamic nature of the AM module improves performance by 1.5, 1.0, and 0.7 mAP over FT set at 0.9, 0.8, and 0.7, respectively. While the randomness of SR and the gradient-only approach in DR limit the performance to 28.7 and 28.3 mAP, the ARR mechanism achieves the performance of 29.2 mAP. These results highlight the significance of the adaptability of AM and ARR over existing methods in dynamic environments. Additional ablation studies on hyper-parameters are provided in the appendix.
5. CONCLUSION
In this work, we proposed AMROD to address the two challenges in Continual Test-Time Adaptation (CTTA). Firstly, object-level contrastive learning leverages ROI features for contrastive learning to refine the feature representation, tailored for object detection. Secondly, the adaptive monitoring module enables efficient adaptation and high-quality pseudo-labels by dynamically skipping unstable adaptation and updating the category-specific threshold, based on the predicted confidence scores. Lastly, the adaptive randomized restoration mechanism selectively reset inactive parameters to mitigate forgetting while retaining valuable knowledge. The empirical results of four CTTA tasks for object detection demonstrate the efficiency and efficacy of AMROD in both short-term and long-term adaptation.
Appendix A Appendix Overview
We provide additional details and analysis in this appendix. In section B, we provide implementation details on AMROD and other baseline methodologies. In section C, we furnish complete results for the quantitative research, additional visualization examples, and additional ablation studies on hyper-parameters. In section D, we further discuss the continual learning. In section E, we discuss the limitations and prospective directions of our research.
Appendix B Additional Implementation Details
Following (VS et al., 2023b), we trained a Faster R-CNN (Ren et al., 2015) with ResNet50 (He et al., 2016) pre-trained on ImageNet (Krizhevsky et al., 2012) on the training set of Cityscapes (Cordts et al., 2016) as the source-trained model of Cityscapes. We utilized an SGD optimizer with a learning rate of 0.001, a momentum of 0.9, and a batch size of 1 and 70k iterations. Similarly, we trained another source model on the clear condition of the SHIFT (Sun et al., 2022) with 30 epochs. We will specify the implementation details of the AMROD and other baseline methods in the following subsections.
B.1. Implememtation of AMROD
Detailed hyper-parameters of AMROD for each benchmark are listed in Table 7. Furthermore, we present a pseudo-code outline of our AMROD framework in Algorithm 1. The three key components of AMROD are highlighted, i.e., Object-level Contrastive Learning, Adaptive Monitoring, and Adaptive Randomized Restoration.
B.2. Implememtation of Baseline Methods
The CoTTA (Wang et al., 2022) and SVDP (Yang et al., 2024) are originally designed for the classification or segmentation tasks. We implement them based on their public code, with some modifications to fit into the objection detection. Specifically, we implement the Cotta method by constructing a mean-teacher framework based on Faster R-CNN, which incorporates a fixed threshold for pseudo-labeling and a stochastic restoration mechanism. Similarly, we implement SVDP with visual prompts on the area of the bounding box in pseudo-labels based on the mean-teacher framework. Further, the MemCLR (VS et al., 2023b) and IRG (VS et al., 2023a) are source-free domain adaptation methods for object detection. Therefore, we implement the MemCLR and IRG based on their public code, with little modifications to follow the CTTA setting (i.e., updating the model parameters during evaluation). For a fair comparison, we set the same momentum $\alpha$ of the Exponential Moving Average (EMA) for CoTTA, SVDP, MemCLR, and IRG as AMROD, all of which apply the mean-teacher framework (Tarvainen and Valpola, 2017). Other hyper-parameters are retained as specified in their public code.
Appendix C Additional Experiment Results
C.1. Quantitative Results
We present the complete experiment results on the two long-term CTTA tasks, i.e., Cityscapes-to-Cityscapes-C and Cityscapes-to-ACDC adaptation tasks, in Table 8 and 9. Experiments show that the proposed AMROD can largely maintain strong performance in the long term.
C.2. Qualitative Results
As shown in figure 4 - 8, we present more visualization results to compare the AMROD and other baseline methods based on the final round adaption on the long-term Cityscapes-to-Cityscapes-C task. The results further validate the effectiveness of AMROD in distinguishing more foreground object categories and better locating them, compared with other baseline methods.
C.3. Additional Ablation Study
To investigate the impact of hyperparameters, we conducted an ablation study on four key values on the long-term Cityscapes-to-Cityscapes-C task, including $\beta_{s}=0.65,0.85$, $\delta_{s}=1.3,1.5$, $\beta_{t}=0.85,0.75$, and $\eta=0.002,0.0005$. The chosen hyper-parameter configuration in this paper of $\beta_{s}=0.75$, $\delta_{s}=1.4$, $\beta_{t}=0.85$, and $\eta=0.001$ yielded a performance of 29.2 mAP. The ablation results, corresponding to different configurations, produced the following mAP scores: 28.8, 29.2, 29.0, 28.9, 29.1, 29.0, and 28.9, respectively. These findings highlight that the selected hyperparameters offer a balanced performance and confirm the stability of CTAOD, as it demonstrates minimal sensitivity to hyperparameter variations.
@@ Appendix E Limitation
Firstly, our experiments are primarily based on a single detector backbone, i.e. Faster R-CNN (Ren et al., 2015). Future work will extend to other advanced detector backbone networks to assess the universality and generalization capabilities of AMROD. Moreover, the proposed ARR mechanism to reset the model every iteration increases the computational cost. Future work could explore more efficient mechanisms for restoration, such as restoring the model only when significant changes in the distribution of the target domain are detected. Finally, the evaluation tasks in our work are designed to simulate real-world adaptation scenarios by incorporating datasets affected by corruption and adverse conditions. However, real-world data distributions are inherently more complex. Consequently, a promising avenue for future research would be to apply our methodology in practical, real-world systems to further validate its effectiveness.
@@
26:        // 3. Adaptive Randomized Restoration
@@
(f) AMROD (ours)
Figure 4. Complete qualitative results. We compare the detection results of the AMROD and other baseline methods in the $10^{th}$ round of adaption to Brightness corruption on the long-term Cityscapes-to-Cityscapes-C task.
@@
(f) AMROD (ours)
Figure 5. Complete qualitative results. We compare the detection results of the AMROD and other baseline methods in the $10^{th}$ round of adaption to Defocus corruption on the long-term Cityscapes-to-Cityscapes-C task.
@@
(f) AMROD (ours)
Figure 6. Complete qualitative results. We compare the detection results of the AMROD and other baseline methods in the $10^{th}$ round of adaption to Fog corruption on the long-term Cityscapes-to-Cityscapes-C task.
@@
(f) AMROD (ours)
Figure 7. Complete qualitative results. We compare the detection results of the AMROD and other baseline methods in the $10^{th}$ round of adaption to Motion corruption on the long-term Cityscapes-to-Cityscapes-C task.
@@
(f) AMROD (ours)
Figure 8. Complete qualitative results. We compare the detection results of the AMROD and other baseline methods in the $10^{th}$ round of adaption to Snow corruption on the long-term Cityscapes-to-Cityscapes-C task.